# OSINT Clinic: Co-designing AI-Augmented Collaborative OSINT Investigations for Vulnerability Assessment

### v2

@@ 7.1.2. Extending the method for collaboration and learning
7.2. Recommendations for streamlining collaborative OSINT investigations
7.2.1. Summary of co-designed ideas
OSINT investigators work with limited resources, in terms of funds, time, and training or software skills to improve their workflows (Glassman and Kang, 2012; Pastor-Galindo et al., 2020). In our work, we bridge the translational gap in HCI research by describing how to leverage generative AI in OSINT-based vulnerability assessments. The learning goals (Table 2) reflect desired outcomes and the design ideas for achieving them can be useful for OSINT practitioners. The challenges from the perspective of students will resonate with novices in the OSINT community, which has been characterized as a community of practice (Belghith et al., 2022). Prior work has described multiple roles played by ChatGPT in the OSINT workflows of experts: (1) research assistance, (2) efficiencies, and (3) technical enablement (Poulter, 2023; Flashpoint, 2024). Though these findings are relevant, our characterization is based on the particular needs of amateurs as they learn OSINT skills, use generative AI to support their collaborative investigations, and deliver high-quality results to clients. Overall, generative AI was useful in training and skill enhancement, scaffolding investigations, and reporting on findings.
Training and skill enhancement
During OSINT investigations, participants used generative AI to understand ethical guidelines and create sock puppet accounts (LG1). They overcame skill gaps in data collection through guidance on effective tools, such as Google dorking and identifying breached data sources (LG3). ChatGPT also supported verification efforts by suggesting alternate sources to participants for cross-referencing information (LG6). AI-assisted interpretation of tool outputs enabled participants to better identify vulnerabilities, enhancing their analytical skills (LG7). Similar benefits like getting new insights and unstuck (Tholander and Jonsson, 2023) have been demonstrated in applications of group ideation (He et al., 2024a), coding (Peng et al., 2023), and general creativity (Doshi and Hauser, 2023).
Scaffolding investigations
ChatGPT helped teams get started and stay organized. It helped participants break down complex OSINT tasks into specific subtasks (LG2) and divide them based on preferences. ChatGPT-generated templates were used to coordinate data collection efforts, reducing redundant work and enhancing communication among team members (LG4). These results contribute to research on using AI to structure complex tasks (Bhattacharjee et al., 2023; Subramonyam et al., 2023).
Reporting on findings
Participants used AI probes to structure raw data into tables and lists, making it easier for team members to follow (LG5). The documentation helped present supporting information for findings in the report. AI was also used to generate actionable recommendations based on identified vulnerabilities, tailoring advice to specific client needs (LG8). Participants iteratively refined their results by simulating client interactions to ensure that recommendations were clear and relevant (LG9). These improvements demonstrate the role of generative AI in the articulation work involved in sensemaking tasks (Schmidt, 2013).
7.2.2. Collaboration and leadership on generative AI platforms
We found that Team-GPT supported three different types of collaboration — sharing, functional, and content — first mentioned in Kang et al.’s longitudinal study on collaboration in intelligence analysis (Kang and Stasko, 2014). The shared workspace in Team-GPT made successful use of AI visible and helped avoid duplication. Functional collaboration was enhanced through features like commenting and forking. Content collaboration was supported by the prompt library and saved chats, which helped create a repository of useful prompts that could be adapted to various contexts. These features can be explored further in the context of team-AI collaboration across other domains (Hendriks et al., 2024).
Leadership behaviors like planning and problem-solving were supported by ChatGPT. Leaders frequently used the prompt library on Team-GPT to create templates that guided the team’s work and structured their presentations. Participants emphasized that AI should complement, not replace, direct team interactions to maintain effective collaboration, aligning with findings from previous studies (Kim et al., 2021; He et al., 2024a).
7.3. Design requirements for generative AI-powered systems for supporting OSINT investigations
Previous research has highlighted several challenges in using generative AI, such as misalignments in understanding its capabilities (Subramonyam et al., 2023; Tankelevitch et al., 2024), prompting techniques (Zamfirescu-Pereira et al., 2023; Subramonyam et al., 2023), and evaluating its outputs (Kulkarni et al., 2023; Subramonyam et al., 2023). Chris Poulter’s blog discusses issues with ChatGPT for OSINT investigations, including operational security, data retention, storage, accuracy, and sourcing (Poulter, 2023). Building on this research, we aimed to identify outstanding challenges within an AI-augmented workflow. Some of the challenges mentioned in the results have been alleviated with recent updates in ChatGPT like providing links to websites as references, formatting content into tables, and generating downloadable files. Here, we outline the key difficulties investigators faced in developing skills and conducting real-world investigations and offer implications for human-AI collaboration research to better adapt foundational models to OSINT-specific applications.
7.3.1. Strengthening privacy measures
Participants were cautious about sharing potentially sensitive public information about companies due to unclear data practices on platforms like ChatGPT and Team-GPT. To mitigate these concerns, they enabled private search (ope, 2024) and made general, scenario-based queries without specifics. OSINT experts have suggested that small, fine-tuned models that can run locally are a promising alternative, given their growing capabilities (Gururangan et al., 2020; pri, 2024). Additionally, software systems can enhance security by adding checkpoints before sending information to models via APIs. For instance, we experimented with running a PII scrubber script locally to sanitize documents containing collected data before using them with ChatGPT.
7.3.2. Integrating existing tools for data collection
There is a lack of support for collecting OSINT information with general-purpose models, even though this is the most time-consuming part of the investigation. Agentic workflows could help automate data collection by orchestrating the use of reputable tools (noa, 2024b; Dennis et al., 2023; Flathmann et al., 2021). These workflows can follow a chain-of-thought reasoning process to document the data collection and analysis based on examples of successful investigations (Jin et al., 2022). Based on design guidelines, improving trust in results can be achieved by integrating well-known and reliable tools (Weisz et al., 2024). With trustworthy AI workflows in place, analysts can focus on verifying information and identifying areas that require deeper investigation.
7.3.3. Improving control over model output
Participants faced challenges with the overwhelming amount of information generated by GPT models, making it difficult to identify the most relevant insights. They highlighted the need for greater control over the output, such as filtering options and creating data visualizations. These features relate to the need for human control over AI-infused systems (Shneiderman, 2022; Yin et al., 2024). Models can be fine-tuned and augmented with knowledge bases to provide the required trustworthy and updated references while suggesting tools and creating templates (Gururangan et al., 2020; Lewis et al., 2020).
7.3.4. Developing prompt templates collaboratively
Participants found it challenging to create effective prompts and wanted to learn from others’ successful examples but struggled to track them. They highlighted the difficulty in achieving consistent results due to significant changes in response with small changes in prompts (Jiang et al., 2022). Identifying reusable prompts was crucial for enhancing the efficiency of their work and getting around ”guardrails” imposed by ChatGPT. Maintaining awareness of how AI is being used and helping team members leverage the output correctly is vital for teamwork (Porsdam Mann et al., 2023). To enable better team awareness while using generative AI tools, there needs to be improved documentation of prompts, their context, and sources of any external information used.
7.3.5. Enhancing leadership support
Effective leadership was crucial for managing investigations and breaking down information silos that formed around individual tasks. To facilitate this, tools should support summarizing ongoing work, identifying areas where help is needed like additional verification, and offering solutions (Hepenstal et al., 2021). This could involve prompting other team members or automatically generating suggestions based on OSINT best practices and guidelines. Personalization based on user behavior and expertise is expected to improve collaborative processes such as planning and monitoring progress, by leveraging the unique perspectives of team members (Wang et al., 2021; Bhattacharjee et al., 2023).
@@ 8. Conclusion
This paper introduced the OSINT clinic, which was aimed at providing students with practical experience in conducting cybersecurity vulnerability assessments using publicly available data, addressing the scalability challenge of helping small businesses protect themselves against cyber attacks. Through a three-phase co-design study, we identified key challenges, such as information management, skill gaps, and collaboration inefficiencies. We developed design ideas to address these gaps by integrating generative AI into various stages of the OSINT investigation process, including planning, data collection, processing, analysis, dissemination, and collaboration. The design ideas can benefit novices and experts to leverage foundational models like ChatGPT and collaborative AI platforms like Team-GPT in their workflows. A pilot with three small businesses demonstrated the benefits of using generative AI to streamline OSINT investigations, as clients found value in the resulting vulnerability assessment reports. It also revealed challenges, including privacy concerns and difficulties in monitoring progress. We propose future research on human-AI collaboration for complex sensemaking, informed by our design requirements and recommendations.

### v3

@@ 7.1.2. Extending the method for collaboration and learning
7.2. Role of Generative AI in Learning and Collaboration for Investigations
Studies 1 and 2 helped identify and achieve a set of learning goals (Table 2) for performing vulnerability assessments as a part of the OSINT Clinic. Generative AI, particularly ChatGPT, was instrumental in helping students plan investigations (LG2) and select effective OSINT tools (LG3). For example, AI simplified the process of breaking down complex tasks into structured subtasks tailored to the team’s capabilities. This scaffolding allowed students to navigate the vast array of OSINT techniques without feeling overwhelmed, supporting novice learners in developing systematic approaches to investigations. These results contribute to research on using AI to structure complex tasks (Bhattacharjee et al., 2023; Subramonyam et al., 2023).
In terms of data collection and verification (LG4, LG6), AI-generated templates provided consistent structures for organizing collected data, reducing redundancy, and enhancing collaboration. AI’s ability to suggest alternative sources for verification encouraged students to adopt rigorous cross-referencing practices. However, the AI’s generalization tendencies and outdated information posed challenges, requiring students to manually verify AI-suggested tools and sources. AI-assisted interpretation of tool outputs enabled participants to better identify vulnerabilities, enhancing their analytical skills (LG7). By simplifying technical findings into clear, understandable language, AI helped students translate complex data into actionable insights. This was particularly beneficial in learning how to present findings to non-technical clients (LG9). Similar benefits like getting new insights and support (Tholander and Jonsson, 2023) have been demonstrated in applications of group ideation (He et al., 2024), coding (Peng et al., 2023), and general creativity (Doshi and Hauser, 2023).
We found that Team-GPT supported sharing and content collaboration, which are collaborative behaviors first identified in Kang et al.’s longitudinal study of teams of intelligence analysts (Kang and Stasko, 2014). Teams used the shared workspace to finalize templates and structure workflows, which minimized duplication of effort and streamlined task division. Content collaboration was supported by the prompt library and saved chats, and led to the development of a repository of useful prompts (see Appendix A.1). However, challenges emerged around maintaining consistency in formatting and style across different team members’ outputs. This required students to develop stronger communication and coordination skills, reflecting real-world team dynamics in cybersecurity investigations (Riebe et al., 2023).
While generative AI supported the development of OSINT skills and collaboration, challenges such as privacy concerns, generalization, and inconsistent outputs underscore the need for careful integration. The students’ ability to critically engage with AI, verify its suggestions, and adapt their workflows reflects a key outcome: they not only learned OSINT techniques but also developed meta-skills in evaluating AI outputs and incorporating them responsibly.
7.3. Design Considerations for Using Generative AI in Real-World Investigations
During the real-world investigations conducted in Study 3, generative AI tools like ChatGPT and Team-GPT played a supportive role across multiple phases of the Intelligence Cycle. Students used generative AI to create templates for structuring data collection, generate actionable recommendations for clients, and simplify technical findings into more digestible reports. Leaders in the team relied on AI to break down complex tasks, assign subtasks, and scaffold the investigation process, ensuring the team maintained a structured workflow. Generative AI also supported the verification process by suggesting alternative data sources, enhancing the rigor of the investigations. However, participants used these tools selectively, often choosing to work independently with OSINT tools and personal ChatGPT instances to maintain efficiency.
The results underscored the need for balancing AI automation with human judgment, aligned with findings from previous studies (Kim et al., 2021; He et al., 2024). While generative AI can handle routine tasks like template generation and summarization, more nuanced activities — such as verifying sensitive information or providing context-aware leadership — still require human expertise. These results build on previous research that has highlighted several challenges in using generative AI, such as misalignments in understanding its capabilities (Subramonyam et al., 2023; Tankelevitch et al., 2024), prompting techniques (Zamfirescu-Pereira et al., 2023; Subramonyam et al., 2023), and evaluating its outputs (Kulkarni et al., 2023; Subramonyam et al., 2023). Integrating generative AI into real-world OSINT workflows revealed design considerations related to privacy, workflow integration, and collaboration. We combined participants’ design ideas with recent technological advancements to propose future directions and mitigation strategies across these challenges.
7.3.1. Privacy Challenges
Privacy concerns were a significant limitation in fully leveraging generative AI tools like ChatGPT and Team-GPT. Although clients did not specifically mention concerns about AI tools or privacy, our NDA explicitly ensured that all data would remain within the clinic and not be shared externally. Participants were cautious about feeding sensitive client data into AI platforms due to uncertainty about how the data might be stored or used by these tools. This caution was more pronounced during the later stages of the investigations, particularly when analysis and dissemination involved sensitive and identifiable information. Even though participants attempted to anonymize data or use generalized scenarios, the fear of exposing confidential information remained prevalent. While a Python script was provided for this purpose, the extra setup was a barrier to its adoption.
The reliance on commercially available, cloud-based AI platforms creates a tension between efficiency and privacy. This issue becomes more critical with OSINT clinics as students may be more prone to taking ”shortcuts” or inadvertently breaching confidentiality. This tension highlights the need for privacy-preserving AI frameworks that can support OSINT investigations without risking client confidentiality. For instance, AI models that run locally on secure machines, rather than cloud-based platforms, could allow for sensitive data processing without risking data leakage (Gururangan et al., 2020; pri, 2024). Features that let users toggle privacy modes, disable chat history, or automatically redact sensitive information can help build trust in these tools. Educational guidelines and training for students on responsible AI use can reduce the likelihood of shortcuts that compromise privacy.
7.3.2. Integration with Existing Workflows
Integrating generative AI into the existing workflows of student analysts presented practical challenges. There is a lack of support for collecting OSINT information with general-purpose models, even though this is the most time-consuming part of the investigation. Analysts also faced workflow disruptions due to context-switching between multiple tools and platforms (ChatGPT, Team-GPT, Google Docs, Mattermost, browser, command line tools).
To address this, future design efforts should focus on creating seamless integrations between generative AI tools and commonly used collaboration platforms. Features like embedded AI prompts within investigation tools or plugin-based extensions could streamline workflows and reduce context-switching. Agentic workflows could help automate data collection by orchestrating the use of reputable tools (noa, 2024b; Dennis et al., 2023; Flathmann et al., 2021). Trust in results can be achieved by integrating well-known and reliable OSINT tools (Weisz et al., 2024; Jin et al., 2022).
7.3.3. Collaboration and Leadership Issues
Generative AI tools introduced both opportunities and challenges for collaboration and leadership within the student teams. Team-GPT features facilitated initial brainstorming and task organization but led to fragmented workflows when used excessively. Forking conversations and managing multiple active chats often created silos, reducing the efficiency of team coordination. As a result, participants preferred to use personal ChatGPT instances to maintain focus and avoid cluttered workspaces.
Effective leadership was crucial in mitigating these challenges. Leaders used AI to structure tasks and generate templates, but monitoring team progress remained difficult. Here, AI can assist leaders by suggesting task assignments based on team members’ skills and past performance (Wang et al., 2021; Bhattacharjee et al., 2023). Features that maintain group awareness, for example, real-time dashboards showing team progress, task ownership, and AI interactions could help leaders and members stay synchronized (Porsdam Mann et al., 2023).
Moreover, while AI facilitated some aspects of task-oriented leadership, it lacked the contextual understanding needed for decision-making. This reinforces the importance of human oversight and interpersonal skills in leadership roles, suggesting that AI should be seen as a complement rather than a replacement for leadership behaviors (Shneiderman, 2022; Yin et al., 2024).
7.4. Limitations and Future Work
A key limitation of this study is the small participant size of six undergraduate students, which may limit the generalizability of the results. However, this is consistent with prior HCI and intelligence analysis research, where studies often rely on small groups due to the sensitive nature of the work (Hepenstal et al., 2021; Riebe et al., 2023; Chin Jr et al., 2009; Kang and Stasko, 2014; Hayes and Cappa, 2018; Gerber et al., 2016; Wong and Kodagoda, 2015). This challenge reflects broader issues in accessing cybersecurity professionals for research. As Sundaramurthy et al. found, barriers such as intensive workloads, sensitive work environments, and mistrust of outsiders make it difficult to recruit larger, more diverse groups (Sundaramurthy et al., 2014).
Another limitation is that the findings are primarily reflective of a small group with limited professional experience. While only half of our student participants were computer science majors, the group represented diverse academic backgrounds, including business and national security majors. Despite this limitation, the insights from this group remain valuable. The participants’ status as novice OSINT analysts mirrors the demographics of many cybersecurity clinics and educational programs, where students are the primary workforce. Their challenges, skill gaps, and learning goals offer a critical lens for understanding how generative AI can support early-stage analysts, who represent the future of the OSINT and cybersecurity workforce. Addressing the needs of novices can lead to the development of AI tools that are accessible, educational, and scalable for broader use in training environments and real-world applications.
The current study provides initial findings on the integration of generative AI in OSINT clinics. The developed prompt library, along with examples of generative AI use across the different phases of OSINT investigations, offers a starting point for refining design ideas in future research. Expanding this work to include larger and more varied participant pools—such as students from diverse backgrounds, professional analysts, or interdisciplinary teams – will help assess the broader applicability and effectiveness of these AI-driven approaches. Including students from various disciplines (e.g., social sciences, law, journalism) could provide a more holistic understanding of how different backgrounds influence the adoption and effectiveness of AI tools in OSINT. Additionally, longitudinal studies that follow participants into professional settings could reveal how AI-supported skills translate to real-world practice.
@@ 8. Conclusion
This paper introduced the OSINT clinic, which was aimed at providing students with training and practical experience in conducting cybersecurity vulnerability assessments using publicly available data, addressing the scalability challenge of helping small businesses protect themselves against cyber attacks. Through a three-phase co-design study, we identified key challenges, such as information management, skill gaps, and collaboration inefficiencies. We developed design ideas to address these gaps by integrating generative AI into various stages of the OSINT investigation process, including planning, data collection, processing, analysis, and dissemination. The design ideas can benefit novices and experts to leverage large language models like ChatGPT and collaborative AI platforms like Team-GPT in their workflows. A pilot with three small businesses demonstrated the benefits of using generative AI to streamline OSINT investigations, as clients found value in the resulting vulnerability assessment reports. It also revealed challenges, including privacy concerns and difficulties in monitoring progress. We propose future research on human-AI collaboration for OSINT tasks, informed by our design requirements and recommendations.